# Distillation-Free One-Step Diffusion for Real-World Image Super-Resolution

### v2

@@ 4.3 Ablation Study
Training Data Scaling. We train DFOSD on the LSDIR (Li et al., 2023) combined with the 10K FFHQ (Karras et al., 2024) dataset and our own collected high-quality dataset, respectively. We provide quantitative results in Table 4 and visual comparisons in Fig. 7. Our collected high-quality dataset provides rich priors, enhancing the authenticity and details.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LR Bicubic LSDIR+FFHQ Our Data
Figure 7: Visual comparison of DFOSD trained on LSDIR+10K FFHQ versus our high-quality dataset.
@@
Figure 8: Visual results ($\times$4) of DFOSD with different perceptual losses. The left side shows a comparison of the checkerboard. The right one shows content about some numbers, i.e., ‘24, 26, 28’.
Loss Function NIQE↓ MUSIQ↑ ManIQA↑ ClipIQA↑ MSE 4.4463 65.35 0.5457 0.5833 LPIPS 4.3331 68.42 0.5914 0.6534 EA-LPIPS 4.1958 68.81 0.6077 0.6519 DISTS 4.2018 69.08 0.6223 0.6555 EA-DISTS 3.9255 69.21 0.6402 0.6683
Table 5: Impact of different perceptual loss functions on DFOSD performance.
Discriminator Base Model NIQE↓ MUSIQ↑ ManIQA↑ ClipIQA↑ None N/A 6.9621 62.36 0.5597 0.5833 Vanilla Discriminator SD 2.1-base 6.1392 64.36 0.5666 0.6059 Diffusion-GAN Discriminator SD 2.1-base 4.5183 67.51 0.5800 0.6246 NAD SD 2.1-base 3.9255 69.21 0.6402 0.6683 NAD SDXL 1.0-base 4.0613 69.86 0.6870 0.6731
Table 6: Performance comparison of DFOSD with different discriminators. The best and second best results of each metric are highlighted in red and blue, respectively.
@@
Perceptual Loss. Table 5 presents the impact of different perceptual loss functions, as well as the scenario where only Mean Squared Error (MSE) is applied as the spatial loss. Figure 8 showcases the visual outcomes of these experiments. The results indicate that incorporating perceptual loss is crucial for training SR models, as it facilitates the generation of more realistic details and enhances overall visual quality. Our proposed edge-aware DISTS (EA-DISTS) achieves the best performance across various image quality metrics and visual assessments. As shown in Fig. 8, EA-DISTS excels in producing highly realistic details, demonstrating its advantage in perceptual quality. This highlights the effectiveness of EA-DISTS in accurately restoring image textures and details, thereby significantly improving the visual quality.
Noise-Aware Discriminator (NAD). We evaluate the impact of various discriminator modules on the training of DFOSD, including NAD, vanilla discriminator, diffusion-GAN (Wang et al., 2023b) style discriminator, and training without any discriminator. Both the vanilla and diffusion-GAN style discriminators are initialized with weights from the Stable Diffusion (SD) 2.1-base model (Rombach et al., 2022a), similar to the NAD described in  Section 3.2. The experimental results, detailed in Table 6, indicate that the generator trained with NAD consistently outperform those utilizing other discriminators that are also initialized with SD 2.1-base. Specifically, NAD demonstrates superior capability in effectively guiding the generator, leading to improved image quality. This demonstrates the advantages of NAD in training distillation-free one-step diffusion models.
@@
Additionally, we conduct experiments where the NAD is initialized with weights from the SDXL (Podell et al., 2023) model to further validate the effectiveness of our approach. As shown in the last two rows of Table 6, the NAD initialized with SDXL 1.0-base weights achieves superior performance compared to its counterparts, without requiring any modifications to the generator’s architecture. This suggests that DFOSD can effectively leverage the strengths of more powerful pre-trained models, and enhance the performance of generator without compromising its efficiency.
5 Differences with other One-Step Diffusion SR Models
We further discuss the difference between our DFOSD and representative one-step diffusion image SR methods, SinSR (Wang et al., 2024b) and OSEDiff (Wu et al., 2024a).
Difference with SinSR. First, SinSR requires performing multi-step deterministic sampling during training to obtain noise-image pairs, which greatly increases the training time. DFOSD does not rely on the results generated by the muti-step pre-trained diffusion models. Second, the involvement of a teacher model during training further escalates memory consumption. In contrast, each training iteration of DFOSD takes a lower latency than SinSR.
Difference with OSEDiff. First, OSEDiff leverages Variational Score Distillation (VSD) to optimize generated images, which necessitates the participation of 3 SD UNets during training, resulting in increased memory usage and prolonged training time. In comparison, our DFOSD requires only 1.5 SD UNets, reducing the training model size by at least $50\%$. Second, OSEDiff extracts prompts from LR images with DAPE, and encoding them into conditional input for SD UNet. DFOSD only uses learnable text embedding as the conditional input, which further reduce computational cost.
@@ 6 Conclusion
In this work, we propose DFOSD, a Distillation-Free One-Step Diffusion model, for Real-ISR. Departing from the diffusion distillation strategies commonly employed in previous studies, our approach effectively reduces training overhead. Specifically, we design a noise-aware discriminator (NAD) that capitalizes on the aggregation capabilities of intermediate features from a pre-trained SD UNet. NAD makes it hard for the generator to distinguish reconstruction from real images. Additionally, we propose the edge-aware DISTS (EA-DISTS) perceptual loss, which significantly enhances the texture realism and visual quality of the generated images. Our distillation-free strategy enables DFOSD to outperform pre-trained multi-step diffusion models in terms of visual results. Comprehensive experiments confirm that DFOSD achieves superior performance and substantially improves the realism of the generated images. These advancements highlight the potential of our method for more efficient and effective image restoration tasks.

### v3

@@ 4.3 Ablation Study
Figure 6: Visual results ($\times$4) of D3SR with different perceptual losses. The left side shows a comparison of the checkerboard. The right one shows content about some numbers, i.e., ‘24, 26, 28’.
@@
Perceptual Loss. Table 5 presents the impact of different perceptual loss functions, as well as only mean squared error (MSE) is applied as the spatial loss. Figure 6 showcases the visual outcomes of these experiments. The results indicate that incorporating perceptual loss is crucial for training SR models, as it facilitates the generation of more realistic details and enhances overall visual quality. Our proposed edge-aware DISTS (EA-DISTS) achieves the best performance across various image quality metrics and visual assessments. As shown in Fig. 6, EA-DISTS excels in producing highly realistic details, demonstrating its advantage in perceptual quality. This highlights the effectiveness of EA-DISTS in accurately restoring image textures and details, thereby significantly improving the visual quality.
Table 5: Impact of different perceptual losses on D3SR.
Loss Function LPIPS↓ NIQE↓ MUSIQ↑ ManIQA↑
MSE 0.3626 4.446 65.35 0.5457
LPIPS 0.3190 4.123 66.41 0.6383
EA-LPIPS 0.3173 4.046 67.47 0.6403
DISTS 0.3463 3.800 67.55 0.6406
EA-DISTS 0.3150 3.747 68.69 0.6436
Table 6: Performance comparison of D3SR with different discriminators. The best and second best results of each metric are highlighted in red and blue, respectively.
Discriminator LPIPS↓ NIQE↓ MUSIQ↑ ManIQA↑
None 0.3862 6.962 62.36 0.5597
CNN 0.3402 6.139 64.36 0.5666
Diffusion-GAN 0.3200 4.518 67.51 0.5800
SD 2.1 (ours) 0.3166 3.925 68.08 0.6198
SDXL (ours) 0.3150 3.747 68.69 0.6436
Diffusion Discriminator. We evaluate the impact of various discriminator modules on the training of D3SR, including our used diffusion discriminator, vanilla discriminator (CNN), diffusion-GAN [48] style discriminator, and training without any discriminator. The CNN discriminator operates in the pixel space, while other discriminators operate in the latent space. The experimental results are shown in Table 6. The comparison between the Diffusion-GAN discriminator and the SD 2.1 discriminator indicates that using a pre-trained Stable Diffusion (SD) model as the discriminator outperforms randomly initialized parameters.
@@
Next, we use SDXL 1.0-base as the discriminator while keeping the generator size unchanged to verify the impact of scaling for diffusion discriminators. As shown in the last two rows of Table 6, the performance of the one-step diffusion model trained with SDXL as the discriminator outperforms the model trained with SD 2.1 as the discriminator, without requiring any modifications to the generator’s architecture. This suggests that D3SR can effectively leverage the strengths of more powerful pre-trained models, and enhance the performance of generator without compromising its efficiency. This breaks the upper bound imposed by the muti-step teacher diffusion model, making further performance improvement simpler.
@@ 5 Conclusion
In this work, we propose D3SR, a One-Step Diffusion model for Real-ISR. Unlike previous methods that use diffusion distillation, our method breaks the limitations of the teacher. We propose the edge-aware DISTS (EA-DISTS) perceptual loss, which enhances the texture realism and visual quality of the generated images. Our adversarial training strategy allows D3SR to outperform multi-step diffusion models in visual quality. Experiments show that D3SR achieves superior performance and improves image realism. This highlights its potential for efficient image restoration.